OpenStudy (anonymous):

Jane walked 4/5 mile on Monday. On Tuesday, she walked 1/2 of the distance she walked on Monday. How many miles did Jane walk on Tuesday? 1/5 mile 2/5 mile 1/2 mile 3/4 mile

OpenStudy (abdela25):

she walked a quarter of the distance she walked on monday.... So she has walked 4 times less.... half the distance would be 1/5

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Can you help with another plz

OpenStudy (abdela25):

sorry i meant B which is 2/5 NOT 1/5

OpenStudy (abdela25):

yeah sure what is it?

OpenStudy (anonymous):

Which fraction is less than -7/12 ? - 2/3 - 5/9 - 2/5 - 1/2

OpenStudy (abdela25):

all of these are negative? or are those your marks

OpenStudy (anonymous):

They are all negative

OpenStudy (abdela25):

okay to make this easier lets make them all positive first do you know how to tell the size differnece of fractions

OpenStudy (abdela25):

like if they are smaller than or larger than equal... etc

OpenStudy (anonymous):

Hmm I can't really remember

OpenStudy (abdela25):

okay let me put it this way if you have a pizza with 12 slices and another pizza with 9 slices which is bigger ?

OpenStudy (anonymous):

12

OpenStudy (abdela25):

technically 12 is bigger but think of it this way a pizza with only 9 slices is going to have bigger slices than a pizza of 12 which is bigger now

OpenStudy (anonymous):

9

OpenStudy (abdela25):

exactly. so for example a 4/9 fraction and a 4/12 fraction... the bigger one is 4/9 if you think of the pizza example

OpenStudy (abdela25):

if there was a pizza with 4 slices and originally had 9 slices these 4 slices are bigger than the 4 slices of a pizza that originally had 12 slices do you understand?

OpenStudy (anonymous):

Yes I understand

OpenStudy (abdela25):

okay now to explain it in the negative fractions

OpenStudy (abdela25):

can you tell me what you think the answer is from your choices? don't worry if you are wrong i'm here to help. im in 10th grade taking High math classes

OpenStudy (anonymous):

Ok. Thanks so much!! I am thinking it's negative 5/9

OpenStudy (abdela25):

very nice job but only if they werent negative

OpenStudy (anonymous):

So did I get it right?

OpenStudy (abdela25):

since they are negative it is the opposite just like in decimals and whole numbers the negative is always the opposite of the positive so here the smaller the denominator the bigger the fraction will be in negative

OpenStudy (abdela25):

lets take a look at it again together

OpenStudy (abdela25):

the fraction -2/3 is also equal to -0.66

OpenStudy (anonymous):

Ok

OpenStudy (abdela25):

do you know how i got that

OpenStudy (anonymous):

No. How? Excuse me I'm blonde. That's why I am making D- in Pre Algebra

OpenStudy (abdela25):

hey look its all good i was once a failure in math too and i'm glad to help you p.s. Blondes are awesome

OpenStudy (anonymous):

Thanks!

OpenStudy (abdela25):

but anyway i got it by dividing them -2 divided by 3 = -0.66

OpenStudy (anonymous):

OHHHH ok

OpenStudy (abdela25):

so knowing that...we can divide the other fractions as well to get there decimals

OpenStudy (anonymous):

Oh ok I get it! ☺

OpenStudy (abdela25):

so can you divide the rest of the fractions for me

OpenStudy (anonymous):

Yes I will divide them

OpenStudy (abdela25):

okay and keep in mind that the SMALLER the negative decimal is the BIGGER it will be form -7/12

OpenStudy (anonymous):

\[-5 ÷ 9 = 1.8\]

OpenStudy (abdela25):

it will always be a negative because a positive (9) and a negative (-5) always equal a negative

OpenStudy (anonymous):

\[-2 ÷ 5 = 0.4\]

OpenStudy (anonymous):

Am I doing it right?

OpenStudy (abdela25):

so -5/9 will = -0.55

OpenStudy (abdela25):

yes you are but you are forgetting the negative sign which is very importnat your teacher will tkae a lot of points off for that

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Thanks

OpenStudy (abdela25):

yeah your welcome :)

OpenStudy (abdela25):

do you want me to divide them for you and show you how i got them

OpenStudy (anonymous):

Yes plz :)

OpenStudy (abdela25):

okay

OpenStudy (abdela25):

the original fraction(-7/12) = -0.58 -2/3 = -0.66 -5/9 = -0.55 -2/5 = -0.40 -1/2 = -0.50

OpenStudy (abdela25):

each of these were easy division and making sure the negative sign was part of the division

OpenStudy (anonymous):

Oh ok

OpenStudy (abdela25):

so your question asks which fraction is bigger than -7/12 correct?

OpenStudy (anonymous):

Correct

OpenStudy (abdela25):

okay since they are all negative everything is opposite the bigger the actual number is the smaller it is

OpenStudy (anonymous):

OK

OpenStudy (abdela25):

with that in mind the smaller the actual number is the bigger it will be towards -7/12

OpenStudy (abdela25):

keep in mind this is only for negatives dont confuse your self with positives this rule only applies to negative numbers only

OpenStudy (abdela25):

can you tell me what you think the answer is now?

OpenStudy (anonymous):

I have a good gut feeling it's -2/5

OpenStudy (abdela25):

yes! exactly! i knew i could teach good! lol but thats the answer

OpenStudy (anonymous):

Thanks so so SO much! you have really helped me!!!

OpenStudy (abdela25):

your welcome any more questions

OpenStudy (anonymous):

Yes one more

OpenStudy (abdela25):

feel free to msg me at anytime for any help in any subject

OpenStudy (abdela25):

okay what is it

OpenStudy (anonymous):

I kinda know how to do this one. But some steps I can't remember. -8 ÷ 2 1/2

OpenStudy (anonymous):

You are so helpful. If I need any help in any other subject I will message you. :D

OpenStudy (abdela25):

okay thanks! so what steps confuse you in this problem

OpenStudy (anonymous):

I can't remember how to start it

OpenStudy (abdela25):

okay i will go step by step from the beginning til the end for you

OpenStudy (anonymous):

Ok thanks!

OpenStudy (abdela25):

first i would like to clarify something is it 2 + 1/2 as 2 differnt numbers (which it probably is) or 2 1/2 as a mixed fraction

OpenStudy (abdela25):

which way is it?

OpenStudy (anonymous):

The 1st way???

OpenStudy (abdela25):

okay so it is -8/2+1/2 right

OpenStudy (anonymous):

right

OpenStudy (abdela25):

okay first you put the fractions in 2+1/2 over a common denominator

OpenStudy (abdela25):

which is 2.2+1/2=2x2/2 + 1/2: -8 over 2x2/2

OpenStudy (abdela25):

are you still with me or are you confused

OpenStudy (anonymous):

Yes. It just sunk in

OpenStudy (abdela25):

okay great

OpenStudy (abdela25):

after that you multiply 2x2 which is 4 so now you have -8 over 4/2+1/2

OpenStudy (abdela25):

when you are confused let me know

OpenStudy (anonymous):

ok

OpenStudy (abdela25):

do you have that taken in

OpenStudy (anonymous):

yes

OpenStudy (abdela25):

ok then you add the fractions over a common denominator to a single fraction will look something like this: 4/2 + 1/2 = 4+1/2: -8/5 over 2

OpenStudy (abdela25):

cool?

OpenStudy (anonymous):

Cool.

OpenStudy (anonymous):

I get it!

OpenStudy (abdela25):

okay good then you have to evaluate or simply 4+1=5 which gives us -8/5 over 2

OpenStudy (abdela25):

do you know what i did here

OpenStudy (anonymous):

yeah kinda

OpenStudy (abdela25):

ok what confuses you

OpenStudy (anonymous):

No wait. I got it